# A Particle Algorithm for Mean-Field Variational Inference

### v1

@@ 1 Introduction
Our convergence criteria resemble those from the literature on Langevin Monte Carlo (LMC), see Dalalyan (2017b); Cheng and Bartlett (2018); Durmus et al. (2019); Durmus and Moulines (2019); Balasubramanian et al. (2022). LMC simulates an SDE whose stationary distribution is proportional to $e^{-V(\cdot)}$, where $V$ is a given potential. Non-asymptotic theoretical guarantees for LMC are given under the condition that the potential $V$ is continuously differentiable, strongly convex, and has Lipschitz gradient; see Durmus and Moulines (2019) for convergence guarantees in the $\mathcal{W}_{2}$ distance. We refer the reader to Chewi (2023) for a survey on LMC.
Our algorithm is built on the framework of gradient flows and particle algorithms. Gradient flows and particle algorithms have recently been applied to a wide range of problems in statistics and optimization. We refer the reader to Liu and Wang (2016); Gallouët and Monsaingeon (2017); Chizat and Bach (2018); Lu et al. (2019); Korba et al. (2020); Lu et al. (2023); Kook et al. (2024); Liu et al. (2024); Yan et al. (2024); Yao et al. (2024) and the references therein for recent developments.
@@ 4 Theoretical analysis
We also note that our algorithm, Algorithm 2, provides a direct approximation $q_{\bm{X}_{n}}$ to the MFVI solution $q_{*}$, with its approximation accuracy guaranteed by Theorem 4.1 and Corollary 4.1. Langevin Monte Carlo (LMC), by contrast, tracks the evolution of a single particle. Theoretical results for LMC (Dalalyan, 2017b; Cheng and Bartlett, 2018; Durmus et al., 2019; Durmus and Moulines, 2019) primarily focus on bounding the $\mathcal{W}_{2}$ distance (or other metrics on probability measures) between the distribution of this particle at the $n$-th iteration and the target distribution. In order to approximate the target distribution using the output of LMC, one needs to generate many such particles and use their empirical distribution. One option is to run $T_{0}$ independent replications of LMC, each time collecting the particle $\mathbf{x}_{t,n}\in\mathbb{R}^{m}$ (where $t\in[T_{0}]$) at the $n$-th iteration, and use the empirical measure $T_{0}^{-1}\sum_{t=1}^{T_{0}}\delta_{\mathbf{x}_{t,n}}$ to approximate the target distribution. Note that this approach would suffer from the curse of dimensionality: in high dimensions, the $\mathcal{W}_{2}$ distance between the empirical distribution and the true one is known to converge very slowly as the sample size $T_{0}$ increases (Weed and Bach, 2019).
@@ 7 Discussions
In this paper, we introduce PArticle VI (PAVI), a particle algorithm for mean-field variational inference (MFVI). The algorithm is based on formulating the relevant optimization problem in terms of Fokker-Planck equations and corresponding diffusion processes. We also provide a theoretical analysis of the particle algorithm, which provides a non-asymptotic upper bound on the $2$-Wasserstein distance between the empirical measure of the particles at the $n$th iteration and the solution of the MFVI optimization problem. Several future directions are worth pursuing. One may consider relaxing the regularity assumptions in Theorem 4.1 and adapting PAVI to the setting of parametric MFVI. Another interesting direction is to develop accelerated versions of PAVI using the recipe in the recent work Chen et al. (2023).
@@ Acknowledgement
The authors thank Daniel Lacker for helpful discussions.

### v2

@@ 1 Introduction
Our convergence criteria resemble those from the literature on Langevin Monte Carlo (LMC), see Dalalyan (2017b); Durmus and Moulines (2017); Cheng and Bartlett (2018); Durmus et al. (2019); Durmus and Moulines (2019); Balasubramanian et al. (2022). LMC simulates an SDE whose stationary distribution is proportional to $e^{-V(\cdot)}$, where $V$ is a given potential. Non-asymptotic theoretical guarantees for LMC are given under the condition that the potential $V$ is continuously differentiable, strongly convex, and has Lipschitz gradient; see Durmus and Moulines (2019) for convergence guarantees in the $\mathcal{W}_{2}$ distance. We refer the reader to Chewi (2023) for a survey on LMC.
Our algorithm is built on the framework of gradient flows and particle algorithms. Gradient flows and particle algorithms have recently been applied to a wide range of problems in statistics and optimization. For example, Saeedi et al. (2017) propose a particle algorithm for approximating discrete distributions with the approximating family given by distributions with at most $K$ atoms (where $K$ is fixed). Cheng et al. (2024) propose particle algorithms for distribution approximation using generalized Wasserstein gradient flow of the Kullback-Leibler divergence. Their algorithms do not impose explicit specification of an approximating family $\mathcal{Q}$ and rely on neural networks to estimate certain vector fields used for updating. Their theoretical analysis relies on assumptions on neural network approximation accuracy and does not consider the error arising from particle approximation. We also refer the reader to Liu and Wang (2016); Gallouët and Monsaingeon (2017); Chizat and Bach (2018); Liu et al. (2019); Lu et al. (2019); Korba et al. (2020); Wang and Li (2020, 2022); Lu et al. (2023); Kook et al. (2024); Liu et al. (2024); Yan et al. (2024); Yao et al. (2024) and the references therein for further developments.
@@ 4 Theoretical analysis
We also note that our algorithm, Algorithm 2, provides a direct approximation $q_{\bm{X}_{n}}$ to the MFVI solution $q_{*}$, with its approximation accuracy guaranteed by Theorem 4.1 and Corollary 4.1. Langevin Monte Carlo (LMC), by contrast, tracks the evolution of a single particle. Theoretical results for LMC (Dalalyan, 2017b; Durmus and Moulines, 2017; Cheng and Bartlett, 2018; Durmus et al., 2019; Durmus and Moulines, 2019) primarily focus on bounding the $\mathcal{W}_{2}$ distance (or other metrics on probability measures) between the distribution of this particle at the $n$-th iteration and the target distribution. In order to approximate the target distribution using the output of LMC, one needs to generate many such particles and use their empirical distribution. One option is to run $T_{0}$ independent replications of LMC, each time collecting the particle $\mathbf{x}_{t,n}\in\mathbb{R}^{m}$ (where $t\in[T_{0}]$) at the $n$-th iteration, and use the empirical measure $T_{0}^{-1}\sum_{t=1}^{T_{0}}\delta_{\mathbf{x}_{t,n}}$ to approximate the target distribution. Note that this approach would suffer from the curse of dimensionality: in high dimensions, the $\mathcal{W}_{2}$ distance between the empirical distribution and the true one is known to converge very slowly as the sample size $T_{0}$ increases (Weed and Bach, 2019).
@@ 7 Discussions
In this paper, we introduce PArticle VI (PAVI), a particle algorithm for mean-field variational inference (MFVI). The algorithm is based on formulating the relevant optimization problem in terms of Fokker-Planck equations and corresponding diffusion processes. We also provide a theoretical analysis of the particle algorithm, which provides a non-asymptotic upper bound on the $2$-Wasserstein distance between the empirical measure of the particles at the $n$th iteration and the solution of the MFVI optimization problem. Several future directions are worth pursuing. One may consider relaxing the regularity assumptions in Theorem 4.1 and adapting PAVI to the setting of parametric MFVI. Another interesting direction is to develop accelerated versions of PAVI using the recipe in the recent works Wang and Li (2022); Chen et al. (2023).
@@ Acknowledgement
The authors thank Daniel Lacker for helpful discussions. Du’s research is supported in part by National Science Foundation grants DMS-2309245 and DMS-1937254, and the Department of Energy grant DE-SC0025347. Wang’s research is supported by the National Science Foundation grant DMS-2210907.